OpenStudy (anonymous):

The polynomial x^3 + 5x^2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.)

OpenStudy (anonymous):

A.) height: 19 in, depth 5 in B.) h 21 in, d 5 in C.) h 19 in, d 7 in D.) h 21 in, d 7 in

OpenStudy (anonymous):

@neer2890

OpenStudy (neer2890):

factor 189 and find which makes the equation zero.

OpenStudy (anonymous):

ok I tried it and got C is that correct??

OpenStudy (neer2890):

how did you find it

OpenStudy (anonymous):

my teacher helped with part of it and i tried to finish the rest that's hoe i got C, i just wanted to check me answer

OpenStudy (neer2890):

you know (x+3) is a factor of \[x ^{3}+5x ^{2}-57x-189=0\] {if you want to check then put x=-3 in above equation} now this whole equation can be written as \[(x+3)(x ^{2}+2x-63)=0\]

OpenStudy (neer2890):

we know that (x+3) is the width of box. so we just have to find the value of x from equation \[x ^{2}+2x-63=0\]

OpenStudy (neer2890):

and we will get the other two dimensions

OpenStudy (anonymous):

how do i do that

OpenStudy (neer2890):

\[x ^{2}+2x-63=0\] =(x+9)(x-7)

OpenStudy (neer2890):

you can find roots of any quadratic equation \[ax ^{2}+bx+c=0\] by this formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

ok and how do i plug that in?

OpenStudy (neer2890):

you will get two values of x as x=-9,7 this indicates that (x+9) and (x-7) are factors of \[x ^{2}+2x-63=0\] because if x=9 then x-9=0 and if x=7 then x-7=0

OpenStudy (neer2890):

now , we know that x+3=15 so, x=15-3=12 so width of box= 12 inches

OpenStudy (neer2890):

similarly, length =(x+9)=12+9=21 inches and depth= x-7=12-7=5 inches

OpenStudy (anonymous):

So B.??

OpenStudy (neer2890):

yup...:p

OpenStudy (anonymous):

thank you so much, i have another question can you help with that one too ill post it in a new question

OpenStudy (neer2890):

ok